Lars:

A cylinder is full at 471 cubic centimeters and has a radius of 5 centimeters. It currently contains 314 cubic centimeters of water. What is the difference between the height of the water in the full cylinder and the height when 314 cubic centimeters of water remains in the cylinder? Use 3.14 for pi. Enter your answer in the box.

Lars:

I think it's 297 cm but I don't know for sure

Joe348:

Use this formula : \[v1=\pi/r^2\]

Joe348:

V= 471 R=3.14 R^2=5 What do you get?

Lemondragon555:

lars i think you need a math tutor, seriously

Lars:

Is this multiplication

@joe348 wrote:
V= 471 R=3.14 R^2=5 What do you get?
I'm doing?

Lars:

@lemondragon555 wrote:
lars i think you need a math tutor, seriously
Yeah I probably do haha

Lemondragon555:

@lars wrote:
@lemondragon555 wrote:
lars i think you need a math tutor, seriously
Yeah I probably do haha
DUH

Joe348:

Okay so basically you're dividing but also multiplying kind of.

Lemondragon555:

huh

Joe348:

So it would be : \[471/(3.14\times 5^2)\]

Joe348:

When you solve that what do you get?

Lemondragon555:

2. something

Lars:

@joe348 wrote:
When you solve that what do you get?
2.88?

Joe348:

You should've got 6!

Lemondragon555:

so 5 squared is 10 so 3.14 x 10 =31.4 so 471 divided by 31.4

Lemondragon555:

@lemondragon555 wrote:
so 5 squared is 10 so 3.14 x 10 =31.4 so 471 divided by 31.4
is this right

Joe348:

5 squared does not each 10

Lemondragon555:

ok

Joe348:

When you square something you multiply.

Lemondragon555:

i know

Lemondragon555:

its just a lot

Joe348:

\[5^2= 5 \times 5=25\]

Joe348:

Okay back to the question did you understand how I solved for the first part of the equation? @lars

Lars:

@joe348 wrote:
Okay back to the question did you understand how I solved for the first part of the equation? @lars
Sorta

Lemondragon555:

soo its 25 x 3.14 = x so its 471 divided by x

Lemondragon555:

@lemondragon555 wrote:
soo its 25 x 3.14 = x so its 471 divided by x
is this rite

Joe348:

You're making it more complicated, please just let me finish!😅

Lemondragon555:

what i want to know

Joe348:

@lars wrote:
@joe348 wrote:
Okay back to the question did you understand how I solved for the first part of the equation? @lars
Sorta
Okay so the question is A cylinder is full at 471 cubic centimeters and has a radius of 5 centimeters. It currently contains 314 cubic centimeters of water. And I solved for the first part of the equation and now we are going to do the same exact thing except we are going to be using 314 cubic for V

Lemondragon555:

@arieonna

Arieonna:

yes

Joe348:

So for the second part of the equation you have : \[314/(3.14\times 5^2)\] \[=4\] Are you understanding it?

Lemondragon555:

@lemondragon555 wrote:
@lemondragon555 wrote:
soo its 25 x 3.14 = x so its 471 divided by x
is this rite
is this right

Lemondragon555:

i need to know

Lars:

@joe348 wrote:
So for the second part of the equation you have : \[314/(3.14\times 5^2)\] \[=4\] Are you understanding it?
Yes it's making a bit more sense, I think I understand how you solved this second part

Joe348:

There's no X in this equation but yes 471 divided by 25 times 3.14

Lars:

@joe348 wrote:
There's no X in this equation but yes 471 divided by 25 times 3.14
Are you talking to me or lemon? Lol

Lemondragon555:

im using x becuz i dont know the real number

Lemondragon555:

@lars wrote:
@joe348 wrote:
There's no X in this equation but yes 471 divided by 25 times 3.14
Are you talking to me or lemon? Lol
i said the equation

Joe348:

@lars wrote:
@joe348 wrote:
So for the second part of the equation you have : \[314/(3.14\times 5^2)\] \[=4\] Are you understanding it?
Yes it's making a bit more sense, I think I understand how you solved this second part
Okay good! Now you're trying to find the difference between the two cylinders

Lemondragon555:

im so confused

Joe348:

So you know the first cylinder equals 6 and the second cylinder equals 4 To find the difference between them you are going to subtract 4 from 6 \[6-4=? \] (Make sure to add centimeters to your answer!)

Lars:

@joe348 wrote:
So you know the first cylinder equals 6 and the second cylinder equals 4 To find the difference between them you are going to subtract 4 from 6 \[6-4=? \] (Make sure to add centimeters to your answer!)
6 minus 4 is 2cm

Joe348:

Yes you got it!

Lemondragon555:

what in the world

Lemondragon555:

i no comprehendo

Lars:

@joe348 wrote:
Yes you got it!
Ohh yay thank you for explaining btw :)

Joe348:

What don't you understand?

Joe348:

Ofc, no problem!

Lemondragon555:

@joe348 wrote:
What don't you understand?
everything

Lars:

Should I close the question now or keep it up since lemon doesn't understand?

Joe348:

Well I have to go, but if still want me to explain it, I'll be back tmrw maybe around 6-7pm!

Joe348:

You can close it.

Lars:

@joe348 wrote:
You can close it.
Okay thank you again!!

Lemondragon555:

yeah close it

Joe348:

:)

Lemondragon555:

but im only 6th grade so